andy101:

During the summer​ heat, ​8 mi bridge expands 8 ft in length. Assuming the bulge occurs straight up the​ middle, how high is the​ bulge?

surjithayer:

we shall consider length of half bridge. expansion is also half and=4 ft half the bridge=1/2*8=4 mile=4*1760*3=21120 ft original length=21120 ft length after expansion=21120+4=21124 ft to find the height we use pythagoras theorem \[h=\sqrt {21124^2-21120^2}=\sqrt{(21124+21120)(21124-21120)}\] \[=\sqrt{42244*4}\] \[=\sqrt{10561*4*4}\] \[=4\sqrt{10561}\] \[\approx 411.07 ~ft\] \[\approx 137 yards.\]